A spherical surface completely surrounds a collection of charges. Find the electric flux through the surface if the collection consists of the following charges.
Answer #1

This is an application of Gauss' law. You only consider the net flux pass through the surface.

Case a: Ea = kq1/r^2

Case b. Eb = kq2/r^2

Case c. Net flux = Ea+Eb

where r is the radius of the sphere.

As values are not given plug in the values